# Turning English-centric LLMs Into Polyglots: How Much Multilinguality Is Needed?

### v1

@@ Results
Figure 2: Average helpfulness of chat responses from Llama 2 7b with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
@@
Figure 3: X-QuAD results for Llama 2 7b with incremental multilingual instruction tuning. Results are shown for both multilingual prompting (en:xx) and monolingual prompting (xx:xx). Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
Figure 2 shows the helpfulness scores assigned by GPT-3.5-Turbo for all models for each of the target languages.
For English, we observe a reasonably strong and stable performance across all incremental multilingual instruction tuning settings, indicating that chat performance does not diminish with the increase in non-English training examples. In contrast, for most non-English target languages, performance increases significantly when moving from monolingual to bilingual instruction tuning. Most notably though, performance tends to plateau when instruction tuning with as few as three different languages. This observation holds for languages in both the supervised setting and the zero-shot setting, indicating that three finetuning languages is sufficient to elicit zero-shot cross-lingual transfer on this task. For languages in the extremely low-resource category, performance remains low despite multilingual finetuning. Manual inspection of these outputs reveal that, while these responses generally look convincing at first glance and are sufficient for language identification (see Figure 1), they are mostly nonsensical. This suggests that useful zero-shot cross-lingual transfer abilities of Llama 2 7b are limited to languages seen with higher frequency during pretraining.
@@ 4.2 Extractive Question Answering
Figure 4: X-CSQA results for Llama 2 7b with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
Figure 5: XNLI results for Llama 2 7b with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
In contrast to open-ended questions commonly used to query LLMs in chat settings, extractive question answering requires the model to identify relevant answer spans within longer context passages provided as part of the prompt. This kind of task closely resembles a retrieval augmented generation (RAG) setting, which is a popular method for extending an LLM’s knowledge with additional data not available during training Lewis et al. (2020); Izacard and Grave (2021). To evaluate our models on this task in multiple target languages, we use XQuAD Artetxe et al. (2020). [^6] We report results measured on the validation split of XQuAD since labels for the test split are not publicly available. This provides 1,190 QA pairs that were professionally translated into different languages.
For this task, we consider both a multilingual prompt setting and a monolingual prompt setting. The former presents the task instruction in English and provides the context passage and question in the relevant target language (en:xx), while the latter is entirely in the target language (xx:xx). As a starting point, we borrow the English prompt from Lai et al. (2023a) and manually translate it into each of the target languages considered. Additionally, we include a standardised response prefix as part of the prompt, effectively force-decoding the response “Based on the passage, the answer to the question is”. This allows us to better isolate the relevant answer string in the generative model’s output. A response is considered correct if the model’s answer string matches the ground truth answer after minimal postprocessing. [^7] We find that some postprocessing of model outputs is necessary for certain languages. Specifically, when queried with German and Russian prompts, all models consistently repeated the question before providing the extracted answer. To handle such edge cases, we strip away the question and truncate the system output to a maximum of 50 characters or the first line break, whichever comes first. An example of our prompting strategy for this task and model outputs is shown in Table 2.
@@ Results
Figure 3 shows the performance on each of our target languages in XQuAD given incremental multilingual instruction tuning.
Similar to the results on the chat task (§4.1), we observe no performance degradation on English as the ratio of multilingual instructions increases. For most non-English target languages in the supervised setting, we see that when using the multilingual prompting strategy (en:xx), performance remains relatively uniform, with the exception of Chinese. In contrast, given the monolingual prompting strategy (xx:xx), multilingual instruction tuning can lead to small but noticeable gains in performance over monolingual tuning. Specifically, for German and Chinese, performance improves when moving from monolingual to bilingual instruction tuning and again plateaus rather quickly with as few as three languages. For languages in the extremely low-resource setting (Hindi and Greek), multilingual instruction tuning fails to improve the model’s ability to handle these languages. While performance gains on this task are generally less pronounced than on the open-ended chat task, we note that zero-shot extractive QA is inherently challenging for LLMs tuned on conversational instructions as they tend to generate verbose responses rather than the single word or entity that correctly matches the ground truth.
@@ 4.3 Commonsense Reasoning
Effectively understanding natural language requires some representation of the natural world and how concepts can relate to one another. Commonsense knowledge describes the set of general facts that reflects this. To evaluate how well English-centric models can reason across languages, we use the X-CSQA dataset. [^8] We report results for X-CSQA measured on the validation set of 1,000 questions. This dataset consists of questions paired with multiple choice answers which aim at assessing general, language-agnostic world knowledge involving different types of commonsense reasoning Talmor et al. (2019); Lin et al. (2021). Given a question and a set of five possible answers from A-E, we prompt the model to output the letter corresponding to the answer that best answers the question. Again, we borrow the English prompt template for this task from Lai et al. (2023a) as a starting point and translate it into each target language. An example of the prompt used is given in Table 3. Finally we assess performance with the monolingual prompting strategy (xx:xx).
@@ Results
Figure 4 shows the accuracy on X-CSQA given incremental multilingual instruction tuning. Again, for English, we observe that multilingual instruction tuning does not significantly degrade performance. However, for non-English target languages, incremental multilingual instruction tuning fails to deliver improvements over monolingual tuning. This contrasts with the results on the previous tasks considered.
Figure 6: Chat performance of Llama 2 7b given multilingual instruction tuning using native non-English (Native NE) examples compared to using translated non-English (MT NE) examples. Error bars show a confidence interval of 95%.
@@ 4.4 XNLI
Natural language inference (NLI) is an important skill for LLMs, especially as input and output sequences grow to consist of multiple sentences. Given two sentences, this task aims to recognise a relationship between them either as entailment, contradiction or neutral. To assess an LLM’s ability to solve this task given multilingual instruction tuning, we use XNLI (Conneau et al., 2018) and evaluate performance on the official test split.
For this task, we use the implementation in the lm-evaluation harness Gao et al. (2023) [^9] https://github.com/EleutherAI/lm-evaluation-harness, which uses language-specific prompts for each target language (i.e. monolingual prompting (xx:xx)). Instead of querying the model to generate the desired label, multiple queries are constructed for each test instance (one for each possible label) and scored by the model. The sequence with the highest likelihood is chosen as the model’s answer.
@@ Results
Figure 5 shows the accuracy measured on our target languages from XNLI given incremental multilingual instruction tuning. Strikingly, performance for all target languages remains uniform regardless of the amount of multilinguality used in finetuning. Furthermore, the equal performance between the Guanaco baseline and our Mono and Multi-$i$ finetuned models suggests that scaling up instruction diversity and the number of languages seen during finetuning beyond just six languages is still insufficient for delivering noticeable gains on this task.
@@ 5 Further Analysis
The results on four distinct evaluation tasks show that multilingual instruction tuning benefits open-ended chat in non-English languages most strongly (§4.1). In this section, we focus on this task to investigate the impact of instruction diversity and model scaling.
@@
Figure 7: Chat performance of Llama 2 70b (denoted by triangular points) and Llama 2 7b (semi-transparent bars) with incremental multilingual instruction tuning. Striped bars indicate that the target language is included in the finetuning data. Error bars show a confidence interval of 95%.
Diversity is a key factor for sample efficient instruction tuning Zhou et al. (2023). Since our experiments in §4 make use of native non-English training instances from the OpenAssistant dataset, a potential confounding factor could be that adding more languages during finetuning also introduces more diverse training instructions. To investigate this, we retrain Multi-$i$ models using translated instruction-tuning data from Mono in place of native non-English examples, following the same incremental recipe as described in §3.2. This ensures that the data distribution remains constant as multilinguality increases. As a translation engine, we again use GPT-3.5-Turbo [^10] Since conversational training instances can be quite long, sometimes exceeding the default 4k token context window, we use gpt-3.5-tubo-16k for this translation task. and the prompt template provided in Figure 10.
Figure 6 compares the resulting performance on the open-ended chat task on a subset of our target languages (results for the remaining target languages are provided in Figure 15). Notably, we observe no significant differences between tuning with distinctly native non-English examples compared to those derived via automatic translation. This indicates that the gains attributed to increased multilinguality are not conflated with an increase in the diversity of instructions.
@@ 5.2 Scaling up Model Size
In order to investigate the effect of model scaling, we also repeat our chat experiments using Llama 2 70b as the underlying base model. Figure 7 shows the resulting helpfulness scores assigned by the LLM judge. Most notably, performance on non-English languages seen relatively frequently during pretraining is dramatically improved, often matching that of English. Secondly, we observe that the larger model’s performance on most non-English target languages tends to plateau with just two finetuning languages, unlike the smaller model that required three. Finally, while the performance on languages in the extremely low-resource scenario remains low, it exhibits a substantial relative improvement compared to the 7-billion parameter model. These results indicate that model scaling is extremely beneficial for exploiting the multilingual capabilities in English-centric models. This aligns with the assertion of Shaham et al. (2023) that larger models are more adept at handling multilinguality.
@@ 6 Discussion
Our findings show that multilingual instruction tuning can elicit zero-shot cross-lingual transfer in English-centric models, though its effectiveness on downstream performance varies across tasks. Notably, we observe significant performance gains for open-ended chat and noticeable gains on extractive QA with monolingual prompting. In comparison, for highly structured tasks such as X-CSQA and XNLI that impose a strict constraint on the output space regardless of the input language (e.g., choosing from options ‘A’, ‘B’, ‘C’, etc.), multilingual instruction tuning does not seem to provide substantial benefits. This distinction highlights that multilingual instruction tuning is most beneficial for open-ended generative tasks where there is an implicit agreement between the input query language and the language output by the model.
We posit that our findings align with the superficial alignment hypothesis Zhou et al. (2023). While instruction tuning guides the model towards a desirable ‘subdistribution of formats’ to use, a small amount of multilingual instruction data allows the model to learn a simple mapping between input and output languages. In essence, it helps steer the model towards language-specific ‘subdistributions of formats’ that agree with the language of the input. Therefore, we conclude that multilingual instruction tuning is of most benefit in scenarios where IO language agreement is required.
Most surprisingly, we find that multilingual instruction tuning with just three languages is sufficient to encourage IO language agreement in smaller LLMs, resulting in improved cross-lingual transfer on generative tasks. As model size increases, however, the number languages required decreases from three to two, indicating that larger models can learn this mapping easier. In both model sizes, adding more instruction tuning languages – including the target language itself – provides no significant gains.
Despite benefits from cross-lingual transfer, our results highlight a clear gap in performance between English and non-English target languages on all tasks. This indicates that small-scale multilingual instruction tuning is insufficient to improve the intrinsic multilingual capabilities of English-centric LLMs and that more advanced interventions may be required (e.g. Pfeiffer et al., 2020; ImaniGooghari et al., 2023).
@@ 7 Conclusion
In this paper, we investigated the use of multilingual instruction tuning to elicit multilingual capabilities of English-centric LLMs. Our results show that finetuning with as few as three languages can promote cross-lingual transfer and allow models to better exploit the relatively small amounts of non-English data seen during pretraining. Experiments conducted on four distinct tasks revealed that this can lead to significant performance improvements on open-ended generative tasks that assume input and output language agreement. Furthermore, we observed that these effects are amplified in larger models, requiring only two finetuning languages for successful multilingual instruction tuning. The effectiveness of cross-lingual transfer, which reduces the need for extensive multilingual instruction tuning, is good news for LLM developers. However, future work could explore other methods to encourage smaller models to better leverage the limited non-English data seen during pretraining and whether there are tasks for which language-specific instruction tuning is of higher importance.
8 Limitations
This work focused on detecting and evaluating non-English responses from English-centric LLMs. Our experiments demonstrated that cross-lingual transfer, which can be elicited given small amount of multilingual instruction tuning, can be effective for improvinging the non-English performance of English-centric models. That said, our analysis focused primarily on Indo-European languages, except for Chinese. As a consequence, it remains unclear how well our results generalise to more distant or topologically diverse languages, however, performance gains observed on Chinese suggest that generalisation is more dependent on exposure during pretraining than language similarity.
@@ 8 Limitations
For our experiments and analysis on model training data, we relied on automatic language identification. To this end, we employed the OpenLID model from Burchell et al. (2023). Despite low error rates achieved by this model, language identification is not perfect and can lead to some texts being misidentified. To mitigate the risk of language contamination Blevins and Zettlemoyer (2022) in our controlled experiments, we only included training examples whose language is identified with a confidence threshold $\geq$ 0.8.
@@
In §5.1 we investigated the impact of multilingual diversity versus training example diversity. While our findings revealed that there is no significant difference between these two settings, we note that even when finetuning with the original native non-English examples, task diversity may be inherently limited by design of the data collection. For instance, regardless of the language used, crowdworkers were asked to follow the same set of guidelines [^11] https://projects.laion.ai/Open-Assistant/docs/guides/guidelines when creating the data.

### v2

@@ Results
Figure 3: SARI weighted by IO language agreement for Llama 2 7b given incremental multilingual instruction tuning. Results are shown for both cross-lingual prompting (en:xx) and monolingual prompting (xx:xx). Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
Figure 2 shows the helpfulness scores assigned by the LLM judge on 16 target languages. For English, performance remains uniform across all multilingual instruction tuning settings. For high- and medium-resource non-English target languages, performance increases significantly when moving from monolingual to bilingual instruction tuning (Multi-2). Meanwhile, performance plateaus when training with as few as three languages (Multi-3), with no substantial differences observed between supervised and 0-shot settings. For low-resource languages, performance remains low despite multilingual finetuning. Manual inspection reveals these outputs, while initially convincing and sufficient for language identification (see Figure 1), are mostly nonsensical. These results indicate that instruction tuning with as few as two to three languages is necessary and sufficient to elicit 0-shot cross-lingual generalisation among high- and medium-resource languages on this task, with language exposure during pretraining being the main limitation to this generalisation ability.
4.2 Sentence Simplification
Sentence simplification aims to make complex sentences easier to read and understand. Due to a lack of high-quality supervised training data, automatic sentence simplification remains a challenging task Štajner (2021), and thus stands to benefit from the few- and 0-shot generalisation capabilities of instruction-following LLMs. To assess performance on this task, we use the MultiSim benchmark Ryan et al. (2023), which includes sentence-aligned parallel datasets in multiple languages. Since the individual datasets in MultiSim are taken from distinct sources, the amount of available data varies across languages. We sample 1,371 complex-simple sentence pairs for each language, except for Slovenian, where we use all 939 instances available (see Appendix D for details).
To prompt models for sentence simplification, we repurpose the detailed instructions given to crowdworkers for the creation of the ASSET corpus Alva-Manchego et al. (2020), which has been shown to be effective in few-shot settings Kew et al. (2023). In its original form, this prompt explicitly states that the output should be suitable for “learners of English”. We remove this language specification so that the model is not explicitly instructed on the target language and instead is expected to infer it from the complex input sentence provided. To translate the prompt into each of our considered target languages, we used a free-tier machine translation service. [^9] https://www.deepl.com/translator We also consider both a cross-lingual prompt setting (en:xx) and a monolingual prompt setting (xx:xx). The former presents the task instruction in English and provides the complex source sentence in the relevant target language, while the latter is entirely in the target language. An example of the prompt template is shown in Table 4.
For evaluation, we report SARI Xu et al. (2016) as an indicator of simplicity and adequacy given the reference simplification. To account for IO language agreement, we weight a model’s corpus-level SARI scores by the percentage of IO language agreement it achieved. This penalises model if it consistently produces outputs in a language that does not match the source sentence. For example, if the model generates English outputs given German input sentences at a rate of 50%, the final score is reduced by half.
Figure 4: XQuAD results for Llama 2 7b given incremental multilingual instruction tuning. Results are shown for both cross-lingual prompting (en:xx) and monolingual prompting (xx:xx). Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
Results
Figure 3 shows the performance according to IO-weighted SARI on each of the target languages considered given incremental multilingual instruction tuning. Similar to the results on the single-turn dialogue task, performance remains uniform for English as the ratio of multilingual instructions increases. For all non-English target languages, we observe a large discrepancy between the cross-lingual prompting strategy and the monolingual prompting strategy. While the former tends to improve as the number of languages increases, it fails to match that of the monolingual prompting strategy under all conditions. In contrast, the results for the monolingual prompting strategy closely resemble those of the single-turn dialogue task across all available target languages: performance gains are most pronounced when moving from monolingual to bilingual instruction tuning, and they generally plateau with as few as three instruction tuning languages.
@@ 4.3 Extractive Question Answering
In contrast to open-ended questions commonly used to query LLMs in chat settings, extractive question answering requires the model to identify relevant answer spans within longer context passages provided as part of the prompt. Such a task closely resembles a retrieval augmented generation (RAG) setting, which is a popular method for extending an LLM’s knowledge with additional data not available during training Lewis et al. (2020); Izacard and Grave (2021). To evaluate our models on this task in multiple target languages we use XQuAD Artetxe et al. (2020). [^10] We report results measured on the validation split of XQuAD since labels for the test split are not publicly available. This provides 1,190 QA pairs that were professionally translated into different languages.
As a starting point, we borrow the English prompt from Lai et al. (2023a) and manually translate it into each of the target languages considered. Additionally, we include a standardised response prefix as part of the prompt, effectively force-decoding the response “Based on the passage, the answer to the question is”. This allows us to better isolate the relevant answer string in the generative model’s output. A response is considered correct if the ground truth answer can be matched with the beginning of the model’s response after minimal post-processing. [^11] We find that some post-processing of model outputs is necessary for certain languages. Specifically, when queried with German and Russian prompts, models consistently repeated the question before providing the extracted answer. To handle such cases, we strip away the repeated question and truncate the response to a maximum of 50 characters or the first line break, whichever comes first. Again, we consider both cross-lingual (en:xx) and monolingual (xx:xx) prompting strategies. An example of our prompting strategy for this task and model outputs is shown in Table 5.
@@ Results
Figure 4 shows the performance on each of our target languages in XQuAD given incremental multilingual instruction tuning. Again, we observe consistent performance for English under multilingual instruction tuning settings. Meanwhile, performance on high-resource languages such as German and Chinese exhibits marginal gains when moving from monolingual to bilingual instruction tuning and plateaus with relatively few languages. For low-resource languages (e.g., Hindi and Greek), performance remains consistently low across all settings, indicating that the model’s ability to generalise to these languages is limited by the lack of exposure during pretraining. Comparing the cross-lingual and monolingual prompting strategies, we observe that the latter generally leads to better performance. This agrees with the results from the sentence simplification task (§4.2) and highlights the benefit of prompting under an IO language agreement assumption. While performance gains on this task are generally less pronounced than on the previous tasks considered, we note that 0-shot extractive QA is inherently challenging for LLMs tuned on conversational instructions as they tend to generate verbose responses rather than the single word or entity that correctly matches the ground truth.
@@ 4.4 Commonsense Reasoning
Figure 5: X-CSQA results for Llama 2 7b given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
Figure 6: XNLI results for Llama 2 7b given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
‘Commonsense knowledge’ is the term frequently used to refer to the set of general facts that reflect how concepts can relate to one another in the real world Li et al. (2022); Talmor et al. (2019). Effectively understanding natural language requires some representation of these concepts and relationships, making commonsense reasoning a key skill for LLMs. To evaluate how well English-centric models can apply commonsense reasoning across languages, we use the X-CSQA dataset Lin et al. (2021). [^12] We report results for X-CSQA measured on the validation set of 1,000 questions. This dataset contains questions paired with multiple choice answers which aim at assessing general, language-agnostic world knowledge involving different types of commonsense reasoning. Given a question and a set of five possible answers from A-E, we prompt the model to output the letter corresponding to the most suitable answer. Again, we borrow the English prompt template for this task from Lai et al. (2023a) as a starting point and translate it into each target language. For this task we consider only the monolingual prompting strategy (xx:xx), which proved most effective in the previous tasks. An example of the prompt used is given in Table 6.
@@ Results
Figure 5 shows the accuracy on X-CSQA given incremental multilingual instruction tuning. Again, for English, we observe that multilingual instruction tuning does not significantly degrade performance. However, in contrast to the previous tasks considered, we do not observe any consistent gains for non-English target languages on the basis of multilingual instruction tuning, suggesting that there are limitations on what tasks this benefits.
Figure 7: Comparison of Llama 2 7b given multilingual instruction tuning using native non-English (Native NE) examples vs. translated non-English (MT NE) examples. Error bars show a 95% confidence interval.
@@ 4.5 XNLI
Natural language inference (NLI) is an important skill for LMs, as it ensures textual coherence, which is particularly critical as input and output sequences grow in length. Given two sentences, this task aims to recognise a relationship between them either as entailment, contradiction, or neutral. To assess an LLM’s ability to solve this task given multilingual instruction tuning, we use XNLI Conneau et al. (2018) and evaluate performance on the official test split.
For this task, we use the implementation in the LM Evaluation Harness Gao et al. (2023). [^13] https://github.com/EleutherAI/lm-evaluation-harness Instead of querying the model to generate the desired label, we use rank classification Sanh et al. (2022); Brown et al. (2020), which allows for direct comparison to the base model. In this setting, multiple queries are constructed for each test instance given a predefined template and subsequently scored by the model. The most probable query sequence under the model is then chosen as the final answer. Prompt templates for this task are language-specific making it equivalent to monolingual prompting (xx:xx). An example of this task is provided in Table 7.
@@ Results
Figure 6 shows the accuracy measured on our target languages from XNLI. Strikingly, performance for all target languages remains uniform regardless of the amount of multilinguality used during finetuning. Furthermore, the comparable performance across all training settings suggests that, in general, instruction tuning fails to provide measurable gains in this challenging task.
@@ 5 Further Analysis
Our experimental results on five distinct tasks demonstrate that multilingual instruction tuning benefits open-ended single-turn dialogue and sentence simplification in non-English languages most strongly. In this section, we focus on single-turn dialogue to investigate the impact of instruction diversity vs. language diversity and model scaling.
@@
Figure 8: Average helpfulness of single-turn dialogue responses from Llama 2 70b (denoted by triangular points) and Llama 2 7b (semi-transparent bars) given incremental multilingual instruction tuning. Striped bars indicate a 0-shot setting and error bars show a confidence interval of 95%.
Diversity is a key factor for sample efficient instruction tuning Zhou et al. (2023). Since our experiments in §4 make use of native non-English training instances from the OpenAssistant dataset, a potential confounding factor could be that adding more languages during finetuning also introduces more diverse training instructions. To investigate this, we retrain Multi-$i$ models using translated instruction tuning data from Mono in place of native non-English examples, following the same incremental recipe as described in §3.2. This ensures that the data distribution remains constant as multilinguality increases. As a translation engine, we again use GPT-3.5-Turbo141414 Since conversational training instances can be quite long, sometimes exceeding the default 4k token context window, we use gpt-3.5-tubo-1106 for this translation task due to its longer context window (16k). and the prompt template provided in Figure 10.
Figure 7 compares the resulting performance on the single-turn dialogue task for a subset of our target languages (results for the remaining target languages are provided in Figure 19). Notably, we observe no significant differences between tuning with distinctly native non-English examples compared to those derived via automatic translation, indicating that gains attributed to increased multilinguality are not conflated with an increase in the diversity of instructions. Following this, in Appendix H, we dive deeper to investigate the role of language diversity given a fixed budget of non-English training examples. From these ablation experiments, we find that the impact of increasing the number of languages is stronger than that of increasing the number of examples for a given language.
@@ 5.2 Scaling up Model Size
In order to investigate the effect of model scaling, we repeat our single-turn dialogue experiments using Llama 2 70b as the underlying base model. Figure 8 shows the resulting helpfulness scores assigned by the LLM judge, with a direct comparison to Llama 2 7b. Most notably, performance on high- and medium-resource non-English target languages is dramatically improved, often matching that of English. However, despite these gains, multilingual instruction tuning with two languages remains essential to elicit cross-lingual generalisation on this task, further underpinning our main findings. In addition, we observe that the larger model’s performance on most non-English target languages tends to plateau with just two finetuning languages, unlike the smaller model that typically required three. Finally, while the performance on low-resource languages remains poor, it exhibits a substantial relative improvement compared to the 7-billion-parameter model. These results indicate that model scaling is extremely beneficial for exploiting the multilingual capabilities in English-centric models and aligns with findings from Shaham et al. (2023).
@@ 6 Discussion
Our findings show that multilingual instruction tuning aids cross-lingual transfer in English-centric LLMs, though its effectiveness on downstream performance varies across tasks and the degree to which a specific target language is seen during pretraining. For high- and medium-resource languages, we observe substantial performance gains on single-turn dialogue and sentence simplification with monolingual prompting, as well as marginal gains on extractive QA with monolingual prompting. Meanwhile, for low-resource languages and highly structured tasks that impose a strict constraint on the output space regardless of the input language (e.g., multiple choice QA), multilingual instruction tuning has little impact. This distinction highlights that multilingual instruction tuning is most beneficial for more open-ended generative tasks that assume IO language agreement, and crucially, that this benefit is limited to high- and medium-resource languages with respect to model’s pretraining language distribution.
Most surprisingly, we find that multilingual instruction tuning with just two to three languages is necessary and sufficient to promote this generalisation in all four of the English-centric LLMs considered. Furthermore, we consistently see that performance tends to plateau rather quickly and adding more instruction tuning languages – including the target language itself – typically provides no significant gains. This underscores that the performance gains are largely the result of cross-lingual transfer and suggests that there may be diminishing returns associated with scaling up multilingual instruction tuning beyond just a handful of languages.
We posit that our findings align with the superficial alignment hypothesis put forward by Zhou et al. (2023), which states that a model acquires its knowledge and abilities during pretraining, while instruction tuning simply guides the model towards a desirable ‘subdistribution of formats’ to use when prompted. We refine this hypothesis by adding that just a small amount of multilingual instruction data also encourages the model to learn a simple mapping between input and output language. This guides the model towards language-specific subdistributions, leading to better performance on tasks where IO language agreement is required.
@@ 7 Conclusion
We investigated the minimal amount of multilingual instruction tuning required to elicit multilingual capabilities of four distinct English-centric LLMs. Our results show that finetuning with as few as two to three languages is necessary and sufficient to promote effective cross-lingual transfer, allowing models to better exploit the relatively small amounts of non-English data seen during pretraining. Experiments on five distinct tasks revealed that this can lead to significant performance improvements for high- and medium-resource languages on open-ended generative tasks that assume input/output language agreement.
While the effectiveness of cross-lingual transfer is indeed good news for LLM developers, future work could explore methods to reduce the performance gap between English and non-English languages, particularly for smaller models, as well as investigating tasks for which language-specific instruction tuning may be of greater importance, such as tasks involving cultural awareness.
Limitations
Our experimental results demonstrate that a relatively small amount of multilingual instruction tuning data can elicit highly valuable cross-lingual transfer, leading to improved performance in open-ended generation tasks that assume IO language agreement. However, we have not explored tasks involving more complicated mappings between languages, such as tasks involving extensive cross-lingual processing or translation.
@@
In this work we rely on automatic language identification to construct multilingual training data and in evaluating model outputs. To this end, we employed the OpenLID model from Burchell et al. (2023). Despite low error rates achieved by this model, language identification is not perfect and can lead to some texts being misidentified. To mitigate the risk of unintentional language contamination Blevins and Zettlemoyer (2022) in our finetuning datasets we include training examples whose language is identified with a confidence threshold $\geq$ 0.8.
@@
In §5.1 we investigate the impact of multilingual diversity versus training example diversity. While our findings reveal that there is no significant difference between these two settings, we note that even when finetuning with the original native non-English examples, task diversity may be inherently limited by design of the data collection. For instance, regardless of the language used, crowdworkers were asked to follow the same set of guidelines [^15] https://projects.laion.ai/Open-Assistant/docs/guides/guidelines when creating the data.
Finally, to evaluate the quality of model-generated responses in the single-turn dialogue task, we rely on LLM-based evaluation methods. While a number of works have shown that LLM-based evaluation provides a decent proxy for assessing the quality of dialogue responses, achieving strong correlations with human judgements Liu et al. (2023); Kocmi and Federmann (2023); Zheng et al. (2023), we stop short of empirically establishing this agreement on our own model outputs. To ensure the validity of model outputs we manually assessed the adequacy of randomly sampled responses in three target languages (en, es, and de). This assessment revealed a similar trend to our LLM-based evaluation, namely that Spanish-language outputs were generally slightly worse than their English-language counterparts but slightly better than those in German.
Ethical Considerations and Risks
This work aims to evaluate multilingual capabilities of English-centric LLMs. In doing so, we acknowledge potential ethical considerations and risks associated with our research. Firstly, LLMs have been shown to inadvertently perpetuate biases present in their training data, which can lead to unexpected and unfair outcomes when these models are used in real-world applications. Therefore, measures must be taken to minimise this risk (e.g., using additional alignment strategies, rigorous testing in multilingual settings) before deploying public-facing models. Secondly, when building on top of English-centric LLMs, there is a risk of cultural homogenization, where nuances and diversity of different languages could potentially be lost due to their under-representation in the training data.